I don't know how to prove that:
2(sin6+cos6)-3(sin4+cos4-1)=0
1
Expert's answer
2010-05-13T09:38:18-0400
Try to use this formulas for your problem:
sin3x = 3sinx - 4(sin^3(x))
cos3x = 4(cos^3(x)) - 3cosx
sin2x = 2sinxcosx
cos2x = cos^2(x) - sin^2(x) = 2(cos^2(x)) - 1 = 1 - 2(sin^2(x))